OpenStudy (anonymous):

solve the system of equations by the elimination method: 11x-8y=92 9x+5y=6 why do you multiplythe first equation by5 and the second by 8

OpenStudy (anonymous):

11x - 8y = 92 (1) 9x + 5y = 6 (2) multiply equation (1) by 5 to get: 55x - 40y = 460 (3) multiply equation (2) by 8 to get: 72x + 40y = 48 (4) now add (3) and (4) to get 127x = 508 now divide by 127 to get x = 508/127 = 4 now use this value of x in eq (1) to find y in summary, we multiply the equations until we can add them together or take them away to eliminate a variable (in this case we did y) then we solve for the remaining variable (x = 4) and use this to solve for y